# FedTSA: A Cluster-based Two-Stage Aggregation Method for Model-heterogeneous Federated Learning

### v1

@@ 0.A.2 Datasets
CIFAR-10 and CIFAR-100. CIFAR-10 and CIFAR-100 are benchmark datasets in the field of Computer Vision (CV), which are also served as popular datasets for evaluating FL methods. CIFAR-10 consists of 60000 32$\times$32 images in 10 classes, with 6000 images per class. CIFAR-100 also consists of 60,000 images but is categorized into 100 classes, each class containing 600 images. CIFAR-10 provides a fundamental platform for image classification tasks, offering a relatively simpler challenge. CIFAR-100, on the other hand, presents a more challenging scenario with more fine-grained categories.
Tiny-ImageNet. Tiny-ImageNet is a simplified version of the larger ImageNet [^4] https://www.image-net.org/ dataset. It consists of 100000 64$\times$64 images in 200 classes, which serves as a bridge between simpler datasets like CIFAR and the full-scale ImageNet dataset. Due to its increased complexity, training on Tiny-ImageNet is more demanding, often used to further assess a method’s efficacy. It is noteworthy that accuracy levels on Tiny-ImageNet tend to be significantly lower compared to those achieved on CIFAR datasets, reflecting its heightened challenge.
@@ 0.A.3 Models and the Pruned Details
In all our experiments, we employ the ResNet18 model, which is also leveraged in baseline studies such as HeteroFL and FedProto. To introduce model heterogeneity, we apply a pruning rate to the hidden size, adjusting the input and output channels. The lower the pruning rate, the simpler the model becomes, whereas the higher the pruning rate, the more complex it is. This pruning process is specifically applied to the first convolutional layer, all subsequent residual blocks, and the input size of the first fully connected layer.
@@ Appendix 0.B Data samples generated by diffusion model
As we can observe, it is evident that the image quality produced by the diffusion model significantly surpasses that of the CGAN-generated images. In Figure 1(a), the majority only exhibit a singular outline, with the corresponding category being hardly recognizable. In contrast, most images in Figure 1(b) accurately depict the corresponding class. High-quality images can facilitate the generation of effective representations by models, thereby enhancing knowledge exchange efficiency and contributing to the overall improvement in model performance.

### v2

@@ 0.A.2 Datasets
CIFAR-10 and CIFAR-100. CIFAR-10 and CIFAR-100 are benchmark datasets in the field of Computer Vision (CV), which are also popular for evaluating FL methods. CIFAR-10 consists of 60,000 32$\times$32 images in 10 classes, with 6000 images per class. CIFAR-100 also consists of 60,000 images but is categorized into 100 classes, each containing 600 images. CIFAR-10 provides a fundamental platform for image classification tasks, offering a relatively simpler challenge. CIFAR-100, on the other hand, presents a more challenging scenario with more fine-grained categories.
Tiny-ImageNet. Tiny-ImageNet is a simplified version of the larger ImageNet [^4] https://www.image-net.org/ dataset. It consists of 100,000 64$\times$64 images in 200 classes, which serves as a bridge between simpler datasets like CIFAR and the full-scale ImageNet dataset. Due to its increased complexity, training on Tiny-ImageNet is more demanding, often used to further assess a method’s efficacy. It is noteworthy that accuracy levels on Tiny-ImageNet tend to be significantly lower compared to those achieved on CIFAR datasets, reflecting its heightened challenge.
@@ 0.A.3 Models and the Pruned Details
In all our experiments, we employ the ResNet18 model, which is also leveraged in baseline studies such as HeteroFL and FedProto. To introduce model heterogeneity, we apply a pruning rate to the hidden size, adjusting the input and output channels. The lower the pruning rate, the simpler the model becomes, whereas the higher the pruning rate, the more complex it is. This pruning process is applied to the first convolutional layer, all subsequent residual blocks, and the input size of the first fully connected layer.
@@ Appendix 0.B Data samples generated by diffusion model
As we can observe, it is evident that the image quality produced by the diffusion model significantly surpasses that of the CGAN-generated images. In Figure 1(a), the majority only exhibit a singular outline, making the corresponding category hardly recognizable. In contrast, most images in Figure 1(b) accurately depict the corresponding class. High-quality images facilitate the generation of effective representations by models, thereby enhancing knowledge exchange efficiency and contributing to the overall improvement in model performance.